# Some one-dimensional elliptic problems with constraints

### v1

@@ 1. Introduction and statement of the results
(1.4) $\mathcal{S}:=\Set{v\in H^{m}(\mathbb{R})}{\int_{\mathbb{R}}v^{2}\,\mathrm{d}x=\rho}$
under suitable conditions on $F$ that include $F^{\prime}(s)=\mathcal{O}(|s|)$ as $s\to 0$. With this approach, $-\lambda$ is nothing but the Lagrange multiplier arising from the constraint $\mathcal{S}$. Since minimisers are among the simplest examples of critical points, it makes sense to wonder whether $J|_{\mathcal{S}}$ is bounded below: this is determined by the behaviour at infinity of $F(s)$ with respect to $|s|^{2+4m}$ and, sometimes, $\rho$. In particular,
$\lim_{|s|\to+\infty}F(s)s^{-(2+4m)}\begin{cases}\leq 0\\ \in(0,+\infty)\\ =+\infty\end{cases}\Longrightarrow\quad\inf_{\mathcal{S}}J>-\infty\begin{cases}\text{for all values of }\rho\\ \text{for small values of }\rho\\ \text{for no values of }\rho\end{cases}$
@@
$-u^{\prime\prime}+\lambda u=F^{\prime}(u)\quad\text{in }\mathbb{R}$
@@
    $F,G\in\mathcal{C}^{1}([0,\infty))$, $F(0)=G(0)=F^{\prime}(0)=G^{\prime}(0)=0$, $G^{\prime}(s)>0$ for all $s>0$, $\lim_{s\to 0^{+}}\frac{F(s)}{G(s)}=0$, and $\lim_{s\to+\infty}\frac{F(s)}{G(s)}=+\infty$.
@@
    $K\in\mathcal{C}([0,\infty))$, $K(0)=0$, $K(s)>0$ for all $s>0$, $s\mapsto\frac{K(s)}{\sqrt{G(s)}}$ is integrable in a right-hand neighbourhood of 00, and $\int_{0}^{\infty}\frac{K(s)}{\sqrt{G(s)}}\,\mathrm{d}s=+\infty$.
@@
$\Phi(t):=\left(\int_{0}^{t}\frac{K(s)}{\sqrt{G(s)}}\,\mathrm{d}s\right)^{2}\ \text{for }t>0,\quad m_{0}:=\max\Set{t\geq 0}{F\leq 0\text{ on }[0,t]},$
$I_{F}:=\sqrt{2}\int_{0}^{m_{0}}\frac{K(s)}{\sqrt{|F(s)|}}\,\mathrm{d}s\in(0,+\infty],\qquad\qquad\qquad$
@@
$L_{0}:=\limsup_{s\to 0^{+}}\frac{Z(s)}{\Phi^{\prime}(s)}\in[0,+\infty],$ $\ell_{0}:=\liminf_{s\to 0^{+}}\frac{Z(s)}{\Phi^{\prime}(s)}\in[0,+\infty],$
@@
Furthermore, $u^{\prime}\in L^{2}(\mathbb{R})$.
@@
    $F\in\mathcal{C}^{1}([0,\infty))$, $\lim_{s\to 0^{+}}\frac{F(s)}{s^{p}}=0$, and $\lim_{s\to+\infty}\frac{F(s)}{s^{p}}=+\infty$.
@@
  • $m_{0}>0$, $I_{F}<\frac{\pi}{\sqrt{(q-p/2+1)K_{\infty}}}$, and $\rho\in\left(\frac{\pi}{\sqrt{(q-p/2+1)k_{\infty}}},I_{F}\right)$;
@@
Furthermore, $u^{\prime}\in L^{2}(\mathbb{R})$.
@@
    The number $62q+2=$ 6, which appears in the definition of $K_{0}$, $k_{0}$, $K_{\infty}$, and $k_{\infty}$, is exactly the exponent (1.5) with $m=1$.
@@
    At first glance, when $F>0$ on $(0,+\infty)$, Corollary 1.2 is weaker than [14, Theorem 1.2] in dimension $1$ (see also [14, Remark 2.6]) because of (a1); however, such an assumption is needed (cf. Remark 2.1 below), hence [14, Theorem 6.6], which [14, Theorem 1.2] is based on, contains a (small) gap.
@@
    Since we can admit sign-changing terms $F$, Corollary 1.2 improves the one-dimensional case of [14, Theorem 1.2]. Additionally, it extends the one-dimensional existence results in [27] to the case of non-linearities more general than the sum of two powers.
@@
$\mathcal{D}:=\Set{v\in H^{m}(\mathbb{R})}{\int_{\mathbb{R}}v^{2}\,\mathrm{d}x\leq\rho},$
which was first used in [4] and then exploited, e.g., in [3, 5, 6, 17, 20, 21, 24, 16]. Moreover, let us recall the Gagliardo–Niremberg inequality [7, 8, 22], here expressed in the one-dimensional case: for every $p>2$ there exists $C_{p}>0$ such that for all $v\in H^{m}(\mathbb{R})$ there holds
@@
If (f0)–(f3) are satisfied and $12\sigma C_{2+4m}^{2+4m}\rho^{2m}<$ 1, then there exist $u\in\mathcal{S}$ and $\lambda>0$ such that $J(u)=\min_{\mathcal{D}}J<0$ and $(\lambda,u)$ is a solution to (1.2).
The proof of Theorem 1.4 follows verbatim that of [24, Theorem 1.1] once proved that every solution $u\in H^{m}(\mathbb{R})$ to the differential equation in (1.2) with $\lambda\in\mathbb{R}$ fixed satisfies the Pohožaev identity if (f0) holds, which is done in Proposition 3.1 below. For this reason, we omit it.
@@
:=$\mathcal{M}:=\Set{v\in H^{m}(\mathbb{R})\setminus\{0\}}{\int_{\mathbb{R}}|v^{(m)}|^{2}\,\mathrm{d}x=\frac{1}{2m}\int_{\mathbb{R}}H(v)\,\mathrm{d}x}.$
@@ 2. The global-branch approach
  • $u_{\lambda}\in\mathcal{C}^{2}(\mathbb{R})$ is a non-negative and even solution of $-u^{\prime\prime}+\lambda G^{\prime}(u)=F^{\prime}(u)$.
@@
with $p>2$ we see that $m_{1}=2\pi$ and $W_{1}^{\prime}(m_{1})=0$, thus no non-trivial non-negative solutions to $-u^{\prime\prime}+u=F^{\prime}(u)$ that vanish at infinity exist in view of [2, Theorem 5].
@@
The function $(0,\infty)\ni\lambda\mapsto m_{\lambda}\in(0,\infty)$ is increasing, continuous, $\lim_{\lambda\to+\infty}m_{\lambda}=+\infty$, and $\lim_{\lambda\to 0^{+}}m_{\lambda}=m_{0}$.
@@
Observe that (A0) implies that $G(s)>0$ if $s>0$. Let $0<\lambda<\Lambda<\infty$ and $s\in(0,m_{\lambda}]$; then,
@@
$0=\lim_{\lambda\to 0^{+}}W_{\lambda}(m_{\lambda})=\lim_{\lambda\to 0^{+}}\lambda G(m_{\lambda})-F(m_{\lambda})=F(M),$
@@
$\rho_{\lambda}:=\int_{\mathbb{R}}K(u_{\lambda}(x))\,\mathrm{d}x=\sqrt{2}\int_{0}^{m_{\lambda}}\frac{K(u)}{\sqrt{G(u)}}\left(\frac{F(m_{\lambda})}{G(m_{\lambda})}-\frac{F(u)}{G(u)}\right)^{-1/2}\,\mathrm{d}u<\infty.$
@@
Since, from (A0) and (A2), $F(s)/G(s)\to 0$ as $s\to 0^{+}$ and $K/\sqrt{G}$ is integrable in a right-hand neighbourhood of 00, we have that $K/\sqrt{W_{\lambda}}$ is integrable in a right-hand neighbourhood of 00 as well. This, together with the property that $W_{\lambda}^{\prime}(m_{\lambda})\neq 0$, yields that $\rho_{\lambda}<+\infty$. Moreover, the continuity of $\lambda\mapsto\rho_{\lambda}$ follows from the one of $\lambda\mapsto m_{\lambda}$ (Lemma 2.2) and the facts above. Finally,
@@
Next, we determine the asymptotic behaviour of $\rho_{\lambda}$ as $\lambda\to 0^{+}$ and $\lambda\to\infty$.
@@
(3) Let us begin with the first part. Given $L\in(L_{\infty},+\infty)$, there exists $S_{L}>m_{0}$ such that $Z(s)\leq L\Phi^{\prime}(s)$ for all $s\geq S_{L}$. In addition, from Lemma 2.2, $m_{\lambda}>S_{L}$ holds for $\lambda\in(\Lambda_{L},\infty)$, with $\Lambda_{L}>0$ large enough. Therefore, we have for $\lambda\in(\Lambda_{L},\infty)$
$\frac{F(m_{\lambda})}{G(m_{\lambda})}-\frac{F(u)}{G(u)}\leq L\Phi(m_{\lambda})-\Phi(u)\quad\text{for all }u\in[S_{L},m_{\lambda}].$
@@
Now we move to the second part. Given $L\in(0,\ell_{\infty})$, there exists $S_{L}>m_{0}$ such that $Z(s)\geq L\Phi^{\prime}(s)$ for all $s\geq S_{L}$. Additionally, since $F/G$ is non-positive on $[0,m_{0}]$ and, from (A1), positive and increasing on $(m_{0},+\infty)$, there holds
@@
In addition, from Lemma 2.2, $m_{\lambda}>S_{L}$ holds for $\lambda\in(\Lambda_{L},\infty)$, with $\Lambda_{L}>0$ large enough. Therefore, we have for $\lambda\in(\Lambda_{L},\infty)$
$\frac{F(m_{\lambda})}{G(m_{\lambda})}-\frac{F(u)}{G(u)}$ $\geq L\Phi(m_{\lambda})-\Phi(u)\quad\text{for all }u\in[S_{L},m_{\lambda}],$
$\frac{F(m_{\lambda})}{G(m_{\lambda})}-\frac{F(u)}{G(u)}$ $\geq L\Phi(m_{\lambda})-\Phi(S_{L})\quad\text{for all }u\in(0,S_{L}],$
@@
$\rho_{\lambda}\leq\sqrt{\frac{2}{L}}\bigl{(}I_{1}(\lambda)+I_{2}(\lambda)\bigl{)}\quad\text{for all }\lambda\in(\Lambda_{L},\infty),$
@@ 3. The poly-harmonic case
$\nabla^{m}u$ :=$:=\begin{cases}\Delta^{m/2}u&\text{if }m\text{ is even,}\\ \nabla\Delta^{(m-1)/2}u&\text{if }m\text{ is odd.}\end{cases}$
@@
Since, from the Sobolev embedding, $|u|_{\infty}<+\infty$, we have that $|g(u)|\lesssim|u|$ a.e. in $\mathbb{R}^{N}$, hence $g\circ u\in L^{2}(\mathbb{R}^{N})\cap L^{\infty}(\mathbb{R}^{N})$. Then, from [26, Lemma 3.1], $u\in W^{2m,p}_{\textup{loc}}(\mathbb{R}^{N})$ for every $p\in[1,+\infty)$. Now, one can follow [5, Proof of Proposition 2.5]. We sketch the proof for the reader’s convenience.
@@
If (F0), (F3), and (1.7) hold, then $\Set{u\in\mathcal{M}}{\varphi_{u}^{\prime\prime}(1)=0}=\emptyset$. As a matter of fact, from $\varphi_{u}^{\prime}(1)=\varphi_{u}^{\prime\prime}(1)=0$ we obtain
@@
We refer to the proofs of [4, Lemma 2.4] or [11, Lemma 2.5 (iv)], which are similar. ∎
We need the following version of Lions’s Lemma (cf. [19, Lemma 3.1]):
@@
$\begin{cases}|G(s)|\leq\varepsilon s^{2}&\text{ if }|s|\in[0,\delta],\\ |G(s)|\leq c_{\varepsilon}|s|^{p}&\text{ if }|s|\in(\delta,M].\end{cases}$
Hence, in view of Lions’ lemma [15, Lemma I.1], we get
@@
Using Lemma 3.8 and proceeding as in [19, Proof of Theorem 1.4], we have the following result in the spirit of [9].
@@
Let $(u_{n})_{n}\subset\mathcal{D}\cap\mathcal{M}$ such that $\lim_{n}J(u_{n})=\inf_{\mathcal{D}\cap\mathcal{M}}J$. From Lemma 3.7, we know that $u_{n}$ is bounded in $H^{m}(\mathbb{R})$. Moreover, Proposition 3.9 applied to the function $H$ implies the existence of the sequences $(\widetilde{u}_{i})_{i=0}^{\infty}\subset H^{m}(\mathbb{R})$ satisfying (3.4) and (3.5). Our first claim is that
@@
$\limsup_{n\to\infty}\frac{1}{2m}\int_{\mathbb{R}}H(u_{n})\,\mathrm{d}x$ $=\limsup_{n\to\infty}\int_{\mathbb{R}}|u_{n}^{(m)}|^{2}\,\mathrm{d}x\geq\sum_{j=0}^{\infty}\int_{\mathbb{R}}|(\widetilde{u}_{j})^{(m)}|^{2}\,\mathrm{d}x$
@@
$=\limsup_{n\to\infty}\frac{1}{2m}\int_{\mathbb{R}}H(u_{n})\,\mathrm{d}x,$
@@
Indeed, given $v\in\mathcal{S}\cap\mathcal{M}$, we consider the curve $(0,\infty)\ni s\mapsto s\star v\in\mathcal{S}$, and the function $\psi_{v}(s)=M(s\star v)=s\varphi^{\prime}_{v}(s)$. We notice that the curve $s\mapsto s\star v$ is not tangent at $v$ to the manifold $\mathcal{M}$, since otherwise we would obtain $\psi^{\prime}_{v}(1)=\varphi^{\prime\prime}_{v}(1)=0$ in contradiction with Remark 3.4). Thus, the manifolds $\mathcal{S}$ and $\mathcal{M}$ do not have the same tangent plane at $v$. Hence, from [20, Proposition A.1], there exist $\lambda\geq 0$ and $\theta\in\mathbb{R}$ such that

### v2

@@ 1. Introduction and statement of the results
[…]
under suitable conditions on $F$ that include $F^{\prime}(s)=\mathcal{O}(|s|)$ as $s\to 0$. With this approach, $\lambda$ is nothing but the Lagrange multiplier arising from the constraint $\mathcal{S}$. Since minimisers are among the simplest examples of critical points, it makes sense to wonder whether $J|_{\mathcal{S}}$ is bounded below: this is determined by the behaviour at infinity of $F(s)$ with respect to $|s|^{2+4m}$ and, sometimes, $\rho$. In particular,
[…]
@@
$u^{\prime\prime}+\lambda u=F^{\prime}(u)\quad\text{in }\mathbb{R}$
@@
    $F,G\in\mathcal{C}^{1}([0,+\infty))$, $F(0)=G(0)=F^{\prime}(0)=G^{\prime}(0)=0$, $G^{\prime}(s)>0$ for all $s>0$, $\lim_{s\to 0^{+}}\frac{F(s)}{G(s)}=0$, and $\lim_{s\to+\infty}\frac{F(s)}{G(s)}=+\infty$.
@@
    $K\in\mathcal{C}([0,\infty))$, $K(0)=0$, $K(s)>0$ for all $s>0$, $s\mapsto\frac{K(s)}{\sqrt{G(s)}}$ is integrable in a right-hand neighbourhood of 0, and $\int_{0}^{+\infty}\frac{K(s)}{\sqrt{G(s)}}\,\mathrm{d}s=+\infty$.
@@
[…]
$I_{F}:=\sqrt{2}\int_{0}^{m_{0}}\frac{K(s)}{\sqrt{|F(s)|}}\,\mathrm{d}s\in(0,+\infty]\text{ (if $m_{0}>0$),}\qquad\qquad$
@@
[…]
@@
Furthermore, $^{\prime}\in L^{2}(\mathbb{R}), uu$ is non-negative and even, and $\lim_{|x|\to+\infty}u(x)=0$.
@@
    $F\in\mathcal{C}^{1}([0,+\infty))$, $\lim_{s\to 0^{+}}\frac{F(s)}{s^{p}}=0$, and $\lim_{s\to+\infty}\frac{F(s)}{s^{p}}=+\infty$.
@@
  • $m_{0}>0$, $I_{F}>\frac{\pi}{\sqrt{(q-p/2+1)k_{\infty}}}$, and $\rho\in\left(\frac{\pi}{\sqrt{(q-p/2+1)k_{\infty}}},I_{F}\right)$;
@@
Furthermore, $^{\prime}\in L^{2}(\mathbb{R}), uu$ is non-negative and even, and $\lim_{|x|\to+\infty}u(x)=0$.
@@
    The number $2q+2=6$, which appears in the definition of $K_{0}$, $k_{0}$, $K_{\infty}$, and $k_{\infty}$, is exactly the exponent (1.5) with $m=1$.
@@
    At first glance, when $F>0$ on $(0,+\infty)$, Corollary 1.2 is weaker than [15, Theorem 1.1] in dimension 11 because of (a1); however, such an assumption is needed (cf. Remark 2.1 below), hence [15, Theorem 6.1], which [15, Theorem 1.1] is based on, contains a (small) gap.
@@
    Since we can admit sign-changing terms $F$, Corollary 1.2 improves the one-dimensional case of [15, Theorem 1.1]. Additionally, it extends the one-dimensional existence results in [28] to the case of non-linearities more general than the sum of two powers.
@@
[…]
which was first used in [4] and then exploited, e.g., in [3, 5, 6, 7, 18, 21, 22, 25, 17]. Moreover, let us recall the Gagliardo–Niremberg inequality [8, 9, 23], here expressed in the one-dimensional case: for every $p>2$ there exists $C_{p}>0$ such that for all $v\in H^{m}(\mathbb{R})$ there holds
@@
If (f0)–(f3) are satisfied and $2\sigma C_{2+4m}^{2+4m}\rho^{2m}<1$, then there exist $u\in\mathcal{S}$ and $\lambda>0$ such that $J(u)=\min_{\mathcal{D}}J<0$ and $(\lambda,u)$ is a solution to (1.2).
The proof of Theorem 1.4 follows verbatim that of [25, Theorem 1.1] once proved that every solution $u\in H^{m}(\mathbb{R})$ to the differential equation in (1.2) with $\lambda\in\mathbb{R}$ fixed satisfies the Pohožaev identity if (f0) holds, which is done in Proposition 3.1 below. For this reason, we omit it.
@@
[…]
@@ 2. The global-branch approach
  • $u_{\lambda}\in\mathcal{C}^{2}(\mathbb{R})$ is a non-negative and even solution of $u^{\prime\prime}+\lambda G^{\prime}(u)=F^{\prime}(u)$.
@@
with $p>2$ we see that $m_{1}=2\pi$ and $W_{1}^{\prime}(m_{1})=0$, thus no non-trivial non-negative solutions to $u^{\prime\prime}+u=F^{\prime}(u)$ that vanish at infinity exist in view of [2, Theorem 5].
@@
The function $(0,+\infty)\ni\lambda\mapsto m_{\lambda}\in(0,+\infty)$ is increasing, continuous, $\lim_{\lambda\to+\infty}m_{\lambda}=+\infty$, and $\lim_{\lambda\to 0^{+}}m_{\lambda}=m_{0}$.
@@
Observe that (A0) implies that $G(s)>0$ if $s>0$. Let $0<\lambda<\Lambda<+\infty$ and $s\in(0,m_{\lambda}]$; then,
@@
$0=\lim_{\lambda\to 0^{+}}W_{\lambda}(m_{\lambda})=\lim_{\lambda\to 0^{+}}\lambda G(m_{\lambda})-F(m_{\lambda})=-F(M),$
@@
$\rho_{\lambda}:=\int_{\mathbb{R}}K(u_{\lambda}(x))\,\mathrm{d}x=\sqrt{2}\int_{0}^{m_{\lambda}}\frac{K(u)}{\sqrt{G(u)}}\left(\frac{F(m_{\lambda})}{G(m_{\lambda})}-\frac{F(u)}{G(u)}\right)^{-1/2}\,\mathrm{d}u<+\infty.$
@@
Since, from (A0) and (A2), $F(s)/G(s)\to 0$ as $s\to 0^{+}$ and $K/\sqrt{G}$ is integrable in a right-hand neighbourhood of 0, we have that $K/\sqrt{W_{\lambda}}$ is integrable in a right-hand neighbourhood of 0 as well. This, together with the property that $W_{\lambda}^{\prime}(m_{\lambda})\neq 0$, yields that $\rho_{\lambda}<+\infty$. Moreover, the continuity of $\lambda\mapsto\rho_{\lambda}$ follows from the one of $\lambda\mapsto m_{\lambda}$ (Lemma 2.2) and the facts above. Finally,
@@
Next, we determine the asymptotic behaviour of $\rho_{\lambda}$ as $\lambda\to 0^{+}$ and $\lambda\to+\infty$.
@@
(3) Let us begin with the first part. Given $L\in(L_{\infty},+\infty)$, there exists $S_{L}>m_{0}$ such that $Z(s)\leq L\Phi^{\prime}(s)$ for all $s\geq S_{L}$. In addition, from Lemma 2.2, $m_{\lambda}>S_{L}$ holds for $\lambda\in(\Lambda_{L},\infty)$, with $\Lambda_{L}>0$ large enough. Therefore, we have for $\lambda\in(\Lambda_{L},+\infty)$
$\frac{F(m_{\lambda})}{G(m_{\lambda})}-\frac{F(u)}{G(u)}\leq L\bigl(\Phi(m_{\lambda})-\Phi(u)\bigr)\quad\text{for all }u\in[S_{L},m_{\lambda}].$
@@
Now we move to the second part. Given $L\in(0,\ell_{\infty})$, there exists $S_{L}>m_{0}$ such that $Z(s)\geq L\Phi^{\prime}(s)$ for all $s\geq S_{L}$. Additionally, since $F/G$ is non-positive on $0,m_{0}]$ and, from (A1), positive and increasing on $(m_{0},+\infty)$, there holds
@@
In addition, from Lemma 2.2, $m_{\lambda}>S_{L}$ holds for $\lambda\in(\Lambda_{L},\infty)$, with $\Lambda_{L}>0$ large enough. Therefore, we have for $\lambda\in(\Lambda_{L},+\infty)$
$\frac{F(m_{\lambda})}{G(m_{\lambda})}-\frac{F(u)}{G(u)}$ $\geq L\bigl(\Phi(m_{\lambda})-\Phi(u)\bigr)\quad\text{for all }u\in[S_{L},m_{\lambda}],$
$\frac{F(m_{\lambda})}{G(m_{\lambda})}-\frac{F(u)}{G(u)}$ $\geq L\bigl(\Phi(m_{\lambda})-\Phi(S_{L})\bigr)\quad\text{for all }u\in(0,S_{L}],$
@@
$\rho_{\lambda}\leq\sqrt{\frac{2}{L}}\bigl(I_{1}(\lambda)+I_{2}(\lambda)\bigl)\quad\text{for all }\lambda\in(\Lambda_{L},+\infty),$
@@ 3. The poly-harmonic case
$\nabla^{m}u$ $:=\begin{cases}\Delta^{m/2}u&\text{if }m\text{ is even,}\\ \nabla\Delta^{(m-1)/2}u&\text{if }m\text{ is odd.}\end{cases}$
@@
Since, from the Sobolev embedding, $|u|_{\infty}<+\infty$, we have that $|g(u)|\lesssim|u|$ a.e. in $\mathbb{R}^{N}$, hence $g\circ u\in L^{2}(\mathbb{R}^{N})\cap L^{\infty}(\mathbb{R}^{N})$. Then, from [27, Lemma 3.1], $u\in W^{2m,p}_{\textup{loc}}(\mathbb{R}^{N})$ for every $p\in[1,+\infty)$. Now, one can follow [5, Proof of Proposition 2.5]. We sketch the proof for the reader’s convenience.
@@
[…]
@@
We refer to the proofs of [4, Lemma 2.4] or [12, Lemma 2.5 (iv)], which are similar. ∎
We need the following version of Lions’s Lemma (cf. [20, Lemma 3.1]):
@@
(3.4) $\begin{cases}|G(s)|\leq\varepsilon s^{2}&\text{ if }|s|\in[0,\delta],\\ |G(s)|\leq c_{\varepsilon}|s|^{p}&\text{ if }|s|\in(\delta,M].\end{cases}$
Hence, in view of Lions’ lemma [16, Lemma I.1], we get
@@
Using Lemma 3.8 and proceeding as in [20, Proof of Theorem 1.4], we have the following result in the spirit of [10].
@@
Let $(u_{n})_{n}\subset\mathcal{D}\cap\mathcal{M}$ such that $\lim_{n}J(u_{n})=\inf_{\mathcal{D}\cap\mathcal{M}}J$. From Lemma 3.7, we know that $u_{n}$ is bounded in $H^{m}(\mathbb{R})$. Moreover, Proposition 3.9 applied to the function $H$ implies the existence of the sequences $(\widetilde{u}_{i})_{i=0}^{+\infty}\subset H^{m}(\mathbb{R})$ satisfying (3.5) and (3.6). Our first claim is that
@@
$\limsup_{n\to+\infty}\frac{1}{2m}\int_{\mathbb{R}}H(u_{n})\,\mathrm{d}x$ $=\limsup_{n\to+\infty}\int_{\mathbb{R}}|u_{n}^{(m)}|^{2}\,\mathrm{d}x\geq\sum_{j=0}^{+\infty}\int_{\mathbb{R}}|(\widetilde{u}_{j})^{(m)}|^{2}\,\mathrm{d}x$
@@
$=\limsup_{n\to+\infty}\frac{1}{2m}\int_{\mathbb{R}}H(u_{n})\,\mathrm{d}x,$
@@
Indeed, given $v\in\mathcal{S}\cap\mathcal{M}$, we consider the curve $(0,+\infty)\ni s\mapsto s\star v\in\mathcal{S}$, and the function $\psi_{v}(s)=M(s\star v)=s\varphi^{\prime}_{v}(s)$. We notice that the curve $s\mapsto s\star v$ is not tangent at $v$ to the manifold $\mathcal{M}$, since otherwise we would obtain $\psi^{\prime}_{v}(1)=\varphi^{\prime\prime}_{v}(1)=0$ in contradiction with Remark 3.4). Thus, the manifolds $\mathcal{S}$ and $\mathcal{M}$ do not have the same tangent plane at $v$. Hence, from [21, Proposition A.1], there exist $\lambda\geq 0$ and $\theta\in\mathbb{R}$ such that